OpenStudy (anonymous):

i want to transpose a function, \[y=243x^5\] is this correct: \[x^5=\frac{y}{243}\Rightarrow x=\sqrt{\frac{y^(1/5)}{3}}\]

OpenStudy (anonymous):

\[x^5=\frac{y}{243}\Rightarrow x=\frac{\sqrt{y^{1/5}}}{3}\]

OpenStudy (anonymous):

so that I can then integrate the area under this curve and y-axis

OpenStudy (anonymous):

i ask, because I keep getting the wrong solution

OpenStudy (anonymous):

\[ \int^{32}_1 \frac{y^{1/5}}{3}=[5\cdot\frac{y^{6/5}}{6\cdot3}]^{32}_1=\frac{5\cdot32\cdot2}{18}=\frac{320}{18}=17.5\]

OpenStudy (anonymous):

However, book solution is 157.5. Can anyone help me here please?

OpenStudy (anonymous):

Sorry, the original question and the follow both contain errors: the follow up should read \[x^5=\frac{y}{243} \Rightarrow x=\frac{y^{1/5}}{3}\]

OpenStudy (anonymous):

@godfreysown can u write the exact question ????????????

OpenStudy (anonymous):

Find area between y=243x^5 and y-axis from y=1 to y=32

OpenStudy (anonymous):

I haven't integrated correctly, that is one thing, but even if one allows for this, my solution is still out by a lot

OpenStudy (anonymous):

I should have subtracted \[\frac{5}{18}\] from 17.5

OpenStudy (anonymous):

but that still gets me no-where near 157.5

OpenStudy (anonymous):

another question: how could I have rewritten the original function so as to translate the curve 45degrees to the left?

OpenStudy (anonymous):

sorry, 90 degrees to the left

OpenStudy (anonymous):

oh, you seem to have given up or lost connection; thanks for beginning to help anyway; cheers

OpenStudy (anonymous):

so\[\int\limits_{1}^{32}[y^(1/5)/3]dy\]  approx :)